# PIR: Remote Sensing Image-Text Retrieval with Prior Instruction Representation Learning

### v1

@@ Abstract
Remote sensing image-text retrieval constitutes a foundational aspect of remote sensing interpretation tasks, facilitating the alignment of vision and language representations. This paper introduces a prior instruction representation (PIR) learning paradigm that draws on prior knowledge to instruct adaptive learning of vision and text representations. Based on PIR, a domain-adapted remote sensing image-text retrieval framework PIR-ITR is designed to address semantic noise issues in vision-language understanding tasks. However, with massive additional data for pre-training the vision-language foundation model, remote sensing image-text retrieval is further developed into an open-domain retrieval task. Continuing with the above, we propose PIR-CLIP, a domain-specific CLIP-based framework for remote sensing image-text retrieval, to address semantic noise in remote sensing vision-language representations and further improve open-domain retrieval performance. In vision representation, Vision Instruction Representation (VIR) based on Spatial-PAE utilizes the prior-guided knowledge of the remote sensing scene recognition by building a belief matrix to select key features for reducing the impact of semantic noise. In text representation, Language Cycle Attention (LCA) based on Temporal-PAE uses the previous time step to cyclically activate the current time step to enhance text representation capability. A cluster-wise Affiliation Loss (AL) is proposed to constrain the inter-classes and to reduce the semantic confusion zones in the common subspace. Comprehensive experiments demonstrate that PIR could enhance vision and text representations and outperform the state-of-the-art methods of closed-domain and open-domain retrieval on two benchmark datasets, RSICD and RSITMD.
@@
Remote Sensing Image-Text Retrieval (RSITR) aims to use image patches (text) to search for matching text (image patches) from large-scale remote sensing databases collected by satellites or aerial drones to obtain valuable information. Due to its important applications [1, 2, 3, 4, 5]in natural resource exploration and disaster monitoring [6, 7], RSITR has received much attention from researchers. In recent years, the rapid expansion of remote sensing data [8, 9] coupled with advancements in cross-modal retrieval techniques [10, 11, 12, 13, 14, 15, 16] has provided substantial data and technical support for RSITR.
@@
We propose a Prior Instruction Representation (PIR) to reduce the influence of semantic noise as shown in Figure 1(b). Firstly, PIR-ITR is designed to utilize prior knowledge of remote sensing scene recognition [38] to perform long-range dependency modelling, ensuring the unbiased vision and text representations, as shown in Figure 2. Two progressive attention encoder (PAE) structures, Spatial-PAE and Temporal-PAE, are proposed to perform long-range dependency modelling to enhance key feature representation. Spatial-PAE utilizes external knowledge to enhance visual-spatial perception. Temporal-PAE is applied to reduce the inaccuracy of position information by leveraging the previous time step to activate the current time step cyclically. In vision representation, a Vision Instruction Representation (VIR) based on Spatial-PAE is proposed to filter features by building a belief matrix, aiming for an unbiased vision representation. In text representation, we design a Language Cycle Attention (LCA) based on Temporal-PAE to obtain an unbiased text representation. For the inter-class similarity [9] in the remote sensing data domain, a cluster-wise attribution loss is proposed to constrain the inter-classes and reduce the semantic confusion zones in the common subspace. Extensive experiments are conducted on two benchmark datasets, RSICD [8], and RSITMD [9], to demonstrate the effectiveness of our approach, resulting in improvements of 4.9% and 4.0% on the RSICD and RSITMD respectively. Based on the PIR, we propose a two-stage CLIP-based method, PIR-CLIP, to address semantic noise in remote sensing vision-language representations and improve open-domain image-text retrieval further. Unlike vanilla CLIP structure, visual prior knowledge is used to instruct visual representations, as shown in Figure 3. It has a universal model architecture and higher retrieval performance than PIR-ITR, which adjusts the visual representation through instruction embeddings and only needs to fine-tune a few parameters, with 7.3% and 9.4% improvement over open-domain state-of-the-art methods on RSICD dataset and RSITMD dataset, respectively.
@@
PIR-ITR is proposed for closed-domain retrieval. Two PAE structures, Spatial-PAE and Temporal-PAE, are proposed to perform long-range dependency modelling to enhance key feature representation. VIR based on Spatial-PAE uses instruction embedding to filter features by building a belief matrix aiming for an unbiased vision representation. LCA based on Temporal-PAE improves text representation by using the previous time step to activate the current time step.
@@ II-A Remote Sensing Image-text Retrieval
Although traditional cross-modal retrieval methods based on natural images are increasingly mature, these methods cannot be directly applied to multiscale and redundant remote sensing images. There is still a long way to go in remote sensing cross-modal retrieval. According to different training paradigms, the remote sensing cross-modal retrieval methods can be divided into two categories: closed-domain and open-domain methods.
@@ II-B Cross-Attention Mechanism
Different from self-attention [44], cross-attention is an attention mechanism in Transformer architecture that mixes two different embedding sequences. These two different embedding sequences come from different modalities or different outputs. Cross-attention [45, 46, 47] is widely used in various deep learning tasks, including image-text retrieval tasks [48, 49, 50, 51].
@@
III METHODOLOGY
Because PIR-ITR is more domain-specific and customized, the methodology introduction first introduces the structure of PIR-ITR. Then, it introduces PIR-CLIP for the structure based on PIR-ITR. This section presents our proposed PIR-ITR (see Figure 2), which introduces the vision encoding and text encoding in Section III-A, progressive attention encoder (PAE) in Section III-B, as a basis for describing Vision Instruction Representation (VIR) and Language Cycle Attention (LCA) in Sections III-C and III-D, respectively, the loss function in Section III-E, training paradigm of PIR-ITR and PIR-CLIP in III-F.
III-A Preliminaries
III-A1 Vision Encoding
Among RSITR methods, common image encoding methods are CNN-based image encoding [20, 42, 21, 22, 41, 52, 9, 23] and Transformer-based image encoding [37]. Since Transformer uses self-attention instead of traditional convolution [53, 54] to make it more powerful for sequence processing, we encode images by using Transformer-based [35] as a vision encoder. An input image is firstly divided into fixed-sized patches. Then these patches are encoded to get a global correlated feature ([CLS]) and local correlated features as
where denotes the vision encoder, is the fine-tuning weights, indicates stacking and concatenation on the sequence length dimension, and is the number of local correlated features.
Some methods [55, 56, 57, 58, 59, 15] introduce some additional prompts or instructions to guide the models for better representations. There are much semantic noise in remote sensing images, which bias the representation of visual semantics. To assist in obtaining unbiased vision representation (see Section III-C), we use ResNet [60] pre-trained on AID dataset [61] as an instruction encoder , where is the pre-trained weights, to obtain instruction embeddings .
III-A2 Text Encoding
@@
In RSITR methods, the RNN-based feature extraction approach [20, 9, 23] and the Transformer-based text feature extraction approach [22, 37] are primary. Compared to RNN [62], Transformer [44] considers all positions in the input sequence simultaneously, thus being able to capture global semantic relationships and long-range dependencies. We use a pre-trained Transformer-based text encoder to encode text . So we can get the global correlated feature ([CLS]) and local correlated features as
@@
where denotes the text encoder, is the fine-tuning weights and is the number of local correlated features.
III-B Progressive Attention Encoder
Existing Transformer-based models [35, 63, 64, 36] are often task-specific and not very flexible to expand according to the task. To enhance the representation of vision and text, two progressive attention encoder (PAE) structures, Spatial-PAE and Temporal-PAE, are designed to perform long-range dependency modeling, which can be used on existing Transformer-based architectures.
III-B1 Transformer Encoder Layer
Transformer Encoder Layer (TEL), as a key component of the Transformer, plays a significant role in natural language processing and computer vision. Our proposed TEL is made up of self-attention [44] and cross-attention [65], as illustrated in Figure 4. Given the packed matrix representations of queries , keys , and values , the scaled dot-product attention is given by
where , and are projection matrices, and is the softmax function. Different from self-attention, cross-attention merges two sequences regardless of modality. Given two different sequences and , the calculation of TEL (see Figure 4) can be expressed as
where is the layer normlization, stands for multilayer perceptron, a common feedforward network. For the purpose of description, we define the TEL calculation as .
III-B2 Messaging between TELs
Based on the way of message transfer between TELs, we designed two different delivery methods, Spatial-PAE and Temporal-PAE, to enhance the key feature representation, as shown in Figure 5.
Spatial-PAE (see Figure 5(a)) makes a spatial connection to the input sequence of the external source using linear projection. This type of connection is not based on fusion but on cross-attention activation. In long-range dependency modeling with TELs, the self-attention mechanism is limited in capturing global information and requires longer sequence distances to establish connections. Spatial-PAE assists long-range dependency modeling with external knowledge that contains global information to enhance the spatial perception of vision and representation of key features.
Temporal-PAE (see Figure 5(b)) makes a temporal connection to the input sequence of the last moment using linear projection. In TEL, positional encoding may not accurately capture the positional relationships within a sequence. When multiple self-attention layers are stacked, potential ambiguity and inaccuracy in the representation of position information can result in performance degradation. Temporal-PAE calculates the attention map using the sequence output from the previous and current time steps, which helps mitigate the ambiguity and inaccuracy of position information with increasing depth of self-attention layers.
III-C Vision Instruction Representation
@@
Unlike natural images, small-scale objects in remote sensing images are more likely to be affected by semantic noise, such as background and irrelevant objects, greatly affecting vision representation. Vision Instruction Representation (VIR) firstly construct the belief matrix using instruction embedding and then filter features for an unbiased vision representation, as shown in Figure 2.
@@
The prior knowledge of remote sensing scene recognition is used to calculate the belief matrix on the local correlated features , and then sort and filter features with a hard belief strategy as
@@
To allow dynamic filtering of features, we propose a soft belief strategy without manually setting the number of features . For batch ranking, the rank of element .
@@
where . This strategy is an alternative to Equation 9 does not require setting the feature size. This strategy makes the retrieval faster and better as shown in Table I. We also use this same strategy in PIR-CLIP to learn unbiased visual representation of large-scale data. The VIR of PIR-CLIP can be specified in Algorithm 2.
Then, Spatial-PAE models long-range dependency relationships on the filtered features activated by external information as
where is the output of -th TEL, and denotes the -th weights of linear projection. Then we can get an unbiased local correlated embedding as
@@
III-D Language Cycle Attention
The text encoder based on Transformer has the ability of bidirectional context modeling, but it still has limitations in a modeling context. Transformer is sensitive to sequence location information [44, 66], which does not explicitly process the position information, but introduces position encoding to the input sequence. This may lead to confusion or ambiguity of position information when dealing with long sequences, resulting in reduced sequence representation capability. Language Cycle Attention (LCA) obtains local correlated text features by cyclically activating word-level features from the previous time step to obtain unbiased text representation.
With Temporal-PAE, we can activate text features recursively as
where is the output of -th TEL, denotes the -th weights of linear projection, and . Then we can get an unbiased local correlated embedding as
@@ III-D Language Cycle Attention
III-E Loss Function
III-E1 Contrastive Loss
In RSITR, the triplet loss [19] is often used to guide the model optimization. Contrastive loss has a better optimization effect compared to the triplet loss (see Table I comparison of (TL) and (CL), so we use contrastive loss [67] to optimize our model, which only requires computing similarities between two samples without additional negative samples.
Given a mini-batch sample, we follow [27, 68] to randomly sample positive remote sensing image-text pairs , with corresponding final vision and text embeddings ( and ) as , and then calculate the cosine similarity as . So we can get the vision-to-text contrastive loss and text-to-vision contrastive loss as
where is the temperature parameter and contrastive loss can be represented as
III-E2 Affiliation Loss
In current, most retrieval tasks only consider pair-wise losses, such as contrastive loss [27, 68] and triplet loss [19], without a comprehensive consideration of the distribution characteristics of the data, as shown in Figure 6(a). We propose affiliation loss, which minimizes the distance between the modality and the clustering center of the affiliated category based on the distribution characteristics of remote sensing data [24], as shown in Figure 6(b). Affiliation loss is a cluster-wise loss [69] to reduce semantic confusion zones [24], based on the contrastive learning and scene category information.
Figure 7 shows the numpy-like pseudocode of affiliation loss. Given a mini-batch sample, we can categorize them into classes based on scene category information 111The RSICD and RSITMD datasets contain category information, which we can get from file name of images. For each image, its corresponding text is grouped into one category, so that we can get the positive pairs , where represents the clustering center corresponding to the -th text. For each text, they are in a similar way as before. We can calculate the cosine similarity of vision-to-text and the cosine similarity of text-to-vision . Then we can obtain the vision-to-text and text-to-vision attribution loss representations as
so affiliation loss can be denoted as

### v2

@@ Abstract
Remote sensing image-text retrieval constitutes a foundational aspect of remote sensing interpretation tasks, facilitating the alignment of vision and language representations. This paper introduces a prior instruction representation (PIR) learning paradigm that draws on prior knowledge to instruct adaptive learning of vision and text representations. Based on PIR, a domain-adapted remote sensing image-text retrieval framework PIR-ITR is designed to address semantic noise issues in vision-language understanding tasks. However, with massive additional data for pre-training the vision-language foundation model, remote sensing image-text retrieval is further developed into an open-domain retrieval task. Continuing with the above, we propose PIR-CLIP, a domain-specific CLIP-based framework for remote sensing image-text retrieval, to address semantic noise in remote sensing vision-language representations and further improve open-domain retrieval performance. In vision representation, we utilize the prior-guided knowledge of the remote sensing scene recognition by building a belief matrix to select key features for reducing the impact of semantic noise. In text representation, we use the previous time step to cyclically activate the current time step to enhance text representation capability. A cluster-wise Affiliation Loss (AL) is proposed to constrain the inter-classes and to reduce the semantic confusion zones in the common subspace. Comprehensive experiments demonstrate that PIR could enhance vision and text representations and outperform the state-of-the-art methods of closed-domain and open-domain retrieval on two benchmark datasets, RSICD and RSITMD.
@@
Remote Sensing Image-Text Retrieval (RSITR) aims to use image patches (text) to search for matching text (image patches) from large-scale remote sensing databases collected by satellites or aerial drones to obtain valuable information. Due to its important applications [1]in natural resource exploration and disaster monitoring [2], RSITR has received much attention from researchers. In recent years, the rapid expansion of remote sensing data [3] coupled with advancements in cross-modal retrieval techniques [4] has provided substantial data and technical support for RSITR.
@@
We propose a Prior Instruction Representation (PIR) to reduce the influence of semantic noise as shown in Figure 1(b). Firstly, PIR-ITR is designed to utilize prior knowledge of remote sensing scene recognition [27] to perform long-range dependency modelling, ensuring the unbiased vision and text representations, as shown in Figure 2. Two progressive attention encoder (PAE) structures, Spatial-PAE and Temporal-PAE, are proposed to perform long-range dependency modelling to enhance key feature representation. Spatial-PAE utilizes external knowledge to enhance visual-spatial perception. Temporal-PAE is applied to reduce the inaccuracy of position information by leveraging the previous time step to activate the current time step cyclically. In vision representation, we filter features by building a belief matrix, aiming for an unbiased vision representation. In text representation, we design Temporal-PAE to obtain an unbiased text representation. For the inter-class similarity [21] in the remote sensing data domain, a cluster-wise attribution loss is proposed to constrain the inter-classes and reduce the semantic confusion zones in the common subspace. Extensive experiments are conducted on two benchmark datasets, RSICD [3], and RSITMD [21], to demonstrate the effectiveness of our approach, resulting in improvements of 4.9% and 4.0% on the RSICD and RSITMD respectively. Based on the PIR, we propose a two-stage CLIP-based method, PIR-CLIP, to address semantic noise in remote sensing vision-language representations and improve open-domain image-text retrieval further. Unlike vanilla CLIP structure, visual prior knowledge is used to instruct visual representations, as shown in Figure 2. It has a universal model architecture and higher retrieval performance than PIR-ITR, which adjusts the visual representation through instruction embeddings and only needs to fine-tune a few parameters, with 7.3% and 9.4% improvement over open-domain state-of-the-art methods on RSICD dataset and RSITMD dataset, respectively.
@@
PIR-ITR is proposed for closed-domain retrieval. Two PAE structures, Spatial-PAE and Temporal-PAE, are proposed to perform long-range dependency modelling to enhance key feature representation. we use instruction embedding to filter features by building a belief matrix aiming for an unbiased vision representation. Temporal-PAE improves text representation by using the previous time step to activate the current time step.
@@ II-B Cross-Attention Mechanism
Different from self-attention [33], cross-attention is an attention mechanism in Transformer architecture that mixes two different embedding sequences. These two different embedding sequences come from different modalities or different outputs. Cross-attention [34] is widely used in various deep learning tasks, including image-text retrieval tasks [35, 36].
@@
III Remote sensing image-text retrieval
We summarise the current remote sensing image-text retrieval (RSITR) methods, first defining a generic image and text encoding scheme and then deforming the presentation of the various available image-text alignment schemes.
III-A Generic Image and Text Encoding
One route to interpreting remote sensing images is aligning images and text at the semantic level. Remote sensing image-text retrieval is the foundation task for this alignment, mainly encoding image and text features through visual and text encoders and then achieving alignment of image and text in the common semantic space. Among RSITR methods, CNN-based image encoder [8, 31, 9, 10, 30, 39, 21, 11] and Transformer-based image encoder [40, 41] are used to image features. Given a RGB-formatted image , it is common practice to encode by a image encoder to get a global correlated feature ([CLS]) and local correlated features as
@@ III-A Generic Image and Text Encoding
where is the fine-tuning weights, indicates stacking and concatenation on the sequence length dimension. If the image encoder is CNN-based, is the high-level features, and is the lower-level feature ensemble. In RSITR methods, the RNN-based feature extraction approach [8, 21, 11] and the Transformer-based text feature extraction approach [10, 26] are primary. Compared to RNN [42], Transformer [33] considers all positions in the input sequence simultaneously, thus capturing global semantic relationships and long-range dependencies. Given a text , it is common practice to get the global correlated feature ([CLS]) and local correlated features as
@@
where denotes the text encoder, is the fine-tuning weights and is the number of local correlated features. For RNN-based text encoders, generally comes from pooled averaging of . Specific methods vary for and extraction.
III-B Various Image-Text Alignment
A simple way to achieve this alignment is by fine-tuning the parameters of vision and text encoders with learnable weights and , aligning and globally, such as the VSE [43]. This method performs better on datasets with a higher proportion of image-text pairs, such as COCO [19], but its improvement is limited in remote sensing applications. Methods such as SCAN [39], CAMP [35] and CAMERA[44] represent images or texts more comprehensively by introducing visual and textual feature-level interactions . These methods do not seem to be able to directly and significantly enhance the remote sensing imagery for representation, as the remote sensing scale is variable, and fine-grained alignment is not easy to achieve. In most RSITR methods, the visual information is expanded from various latitudes because the remote sensing imagery contains many redundant information. AMFMN [21] and SWAN [12] perform multi-scale feature fusion of and and extracts salient features of remote sensing images by redundant feature filtering. GaLR [11] incorporates and external object-level knowledge processed by graph convolutional networks (GCN), with feature-level interactions . These methods are aligned to optimised visual and textual features.
IV Prior Instruction Representation Learning
We first define prior instruction representation (PIR) learning in RSITR and then provide a detailed description of PIR-ITR, CLIP-baed PIR-CLIP.
IV-A Preliminaries
There are much semantic noise in remote sensing images, which bias the representation of visual semantics. Some methods [45, 46, 47, 48, 49, 50] introduce some additional prompts or instructions to guide the models for better representations. Taking inspiration from these methods, we use ResNet [51] pre-trained on AID dataset [52] as an instruction encoder , where is the pre-trained weights, to obtain instruction embeddings . As mentioned above GaLR [11] incorporates and external object-level knowledge processed by GCN, but its role as an element of feature fusion may influence the intrinsic representations of the image. Different from GaLR, instruction embeddings serve as instructions to help visual feature representation. We can express this as a pseudo-formula as . Our goal is to filter out unbiased visual features from with the help of .
IV-B Belief Strategy
@@
Unlike natural images, small-scale objects in remote sensing images are more susceptible to semantic noise from background or irrelevant objects, significantly impacting visual representation. The belief strategy addresses this by first constructing a belief matrix using instruction embeddings, and then filtering features to achieve an unbiased visual representation, as shown in Figure 2.
@@
The prior knowledge of remote sensing scene recognition is used to calculate the belief matrix on the local correlated features , and then sort and filter features with a hard belief strategy [28] as
@@
Despite the fact that the above strategy in practice, an empirical setting of the appropriate number of features is required. To allow dynamic filtering of features, we propose a soft belief strategy without . For batch ranking, the rank of element .
@@
where . The new features are score-based and ranking-based i.e. relative ranked weighted features. This strategy is an alternative to Equation 4 does not require setting the number of visual features. This strategy makes the retrieval faster and better as shown in Table I. We also use this same strategy in PIR-CLIP to learn unbiased visual representation from large-scale data training.
IV-C Image-Text Alignment with PIR Learning
IV-C1 Progressive Attention Encoder Layer
Transformer Encoder Layer (TEL), a crucial element of the Transformer architecture, is integral to natural language processing and computer vision tasks. We design a Progressive Attention Encoder Layer (PAEL) based on TEL, which incorporates self-attention [33] and cross-attention [53], as shown in Figure 3. Different from self-attention, cross-attention merges two sequences regardless of modality. Given two different sequences and , the calculation of PAEL (see Figure 3) can be expressed as
where is multi-head attention, is the layer normlization, stands for multilayer perceptron, a common feedforward network. For the purpose of description, we can define the PAEL calculation as .
IV-C2 Progressive Attention Encoder
We developed two distinct message transfer methods between PAELs: Spatial-PAE (S-PAE) and Temporal-PAE (T-PAE), to capture key feature representation, as illustrated in Figure 4.
Spatial-PAE (see Figure 4(a)) establishes a spatial connection to the input sequence from an external source using linear projection. Unlike fusion-based approaches [12, 54, 11], this connection relies on cross-attention. In long-range dependency modeling with PAELs, the self-attention mechanism has limited capacity to capture global information and requires extended sequence distances to form connections. Spatial-PAE models long-range dependency relationships on the filtered features activated by external information as
where is the output of -th PAEL, , and denotes the -th weights of linear projection. Then we can get an unbiased local correlated embedding as
@@ IV-C2 Progressive Attention Encoder
Temporal-PAE (see Figure 4(b)) establishes a temporal connection to the input sequence of the previous time step via linear projection. In TEL, positional encoding may not effectively capture positional relationships within a sequence. Temporal-PAE computes the attention map using outputs from both the previous and current time steps, addressing the ambiguity and inaccuracies in position information that arise with deeper self-attention layers. This allows Temporal-PAE to recursively activate text features, which can be expressed as follows
where is the output of -th PAEL, , denotes the -th weights of linear projection, and . Then we can get an unbiased local correlated embedding as
@@
IV-C3 PIR-ITR and PIR-CLIP
According to the PIR learning idea, we first designed the soft belief strategy in inheritance of the previous hard belief [28] strategy of PIR-ITR. For PIR-ITR with soft belief strategy, we first obtain positive remote sensing image-text pairs from the remote sensing image-text datasets . Then, the image and text are encoded separately to obtain the corresponding visual features , instruction embeddings and textual features . After that, we compute the belief matrix of visual features and instruction embeddings and calculate the rank of element in batch training. For each element of dimension in , we get rank-weighted visual features . Finally, we can use Spatial-PAE and Temporal-PAE to get final and respectively.
The PIR-ITR directly performs the PIR learning paradigm after visual and text encoder processing for the above closed-domain retrieval. However, this complex and large structure is not easy to train on large-scale data. For PIR-CLIP, we first pre-train on RS5M dataset [17] using CLIP with contrastive loss, then fine-tune visual representation from RSITR datasets. The fine-tuning process is similar to PIR-ITR, with the key difference being that the visual backbone is updated using the base ViT and ResNet models, while the text encoder is replaced by a language transformer trained on a large-scale remote sensing corpus, eliminating the need for an additional Spatial-PAE. Unlike PIR-ITR, PIR-CLIP modifies the ViT structure by employing convolution for patching, followed by a soft belief strategy, and then a transformer encoder to facilitate interaction between visual and instructional features.
IV-D Loss Function
IV-D1 Contrastive Loss
In RSITR, triplet loss [7] is commonly used for model optimization. However, contrastive loss provides better optimization performance, as shown in the comparison between (TL) and (CL) in Table I. Therefore, we adopt contrastive loss [55] to optimize our model, as it requires only pairwise similarity calculations without the need for additional negative samples.
Given the and embedded by the PIR-ITR or PIR-CLIP from batch image-text pairs, we calculate the cosine similarity of them as . Then we can get the contrastive loss as
where is the temperature parameter.
IV-D2 Affiliation Loss
Most current retrieval tasks focus solely on pair-wise losses, such as contrastive loss [15, 56] and triplet loss [7], without fully accounting for the distributional characteristics of the data (see Figure 5(a)). To address this, we propose affiliation loss, which leverages the distributional characteristics of remote sensing data to minimize the distance between each modality and the clustering center of its affiliated category [12], as illustrated in Figure 5(b).
Affiliation loss is a cluster-based loss [57] designed to reduce semantic confusion zones [12] by incorporating contrastive learning and scene category information. Figure 6 presents numpy-like pseudocode for the affiliation loss. Given a mini-batch sample, we can categorize them into classes based on scene category information 111The RSICD and RSITMD datasets contain category information, which we can get from file name of images.. For each image, its corresponding text is grouped into one category, so that we can get the positive pairs , where represents the clustering center corresponding to the -th text. For each text, they are in a similar way as before. We can calculate the cosine similarity of vision-to-text and the cosine similarity of text-to-vision . Then we can obtain the attribution loss representations as